gcf and quadratic equation
4xsquared-8x-60=0

Answer by dawnhannah(7)

Well this question is a quadradic equation because there is an x squared term. To solve you want to factor out the greatest common factor, and that is a 4.
4x^2-8x-60=0 will become 4(x^2-2x-15)=0
Next you would take the x^2-2x-15=0 and factor that. Multiply the first and last coefficients which in this case is 1 and 15. Then find the factors of 15 that add to the middle term of 2. Break the middle term into two terms using those factors as the coefficients.
So broken down it looks like this.
4x^2-8x-60=0
4(x^2-8x-60)=0
4(x-5)(x+3)=0
Lastly you would set them equal to 0.
x-5=0 and x+3=0
And solve for x by adding 5 to both sides
x-5=0
+5 +5
x=5
And subtracting 3 to both sides
x+3=0
-3 -3
x=-3

x=5 and x=-3